# Scale-Invariant Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

### v1

@@ A.2 Proof of Theorems
Let a policy $\pi_{\bar{\theta}}$ is constrained-Pareto optimal. Then the following equation holds:
$\bar{\theta}=\mathrm{argmin}_{\theta}-J_{R_{1}}(\theta)\;\textbf{s.t.}\;J_{R_{i}}(\bar{\theta})\leq J_{R_{i}}(\theta)\;\forall i\geq 2,\;J_{C_{k}}(\theta)\leq d_{k}\;\forall k.$
Since the gradients of the objective and constraint functions are continuous, we can get the following by linear approximation at $\theta=\bar{\theta}$:
$\mathbf{0}=\mathrm{argmin}_{d}-\nabla J_{R_{1}}(\bar{\theta})^{T}d\;\textbf{s.t.}\;0\leq\nabla J_{R_{i}}(\bar{\theta})^{T}d\;\forall i\geq 2,\;J_{C_{k}}(\bar{\theta})+\nabla J_{C_{k}}(\bar{\theta})^{T}d\leq d_{k}\;\forall k.$ (28)
The dual problem is as follows:
$\underset{\lambda\geq 0,\nu\geq 0}{\mathrm{max}}L(\lambda,\nu):=\mathrm{min}_{d}-\nabla J_{R_{1}}(\bar{\theta})^{T}d-\sum_{i=2}^{N}\lambda_{i}J_{R_{i}}(\bar{\theta})^{T}d+\sum_{k=1}^{M}\nu_{k}(J_{C_{k}}(\bar{\theta})-d_{k}+\nabla J_{C_{k}}(\bar{\theta})^{T}d).$
Since linear programming problems hold strong duality, the stationarity of the KKT condition is satisfied at the optimal dual variables $\lambda^{*}$, $\nu^{*}$, and $d=\mathbf{0}$ as follows:
$-\nabla J_{R_{1}}(\bar{\theta})-\sum_{i=2}^{N}\lambda^{*}_{i}J_{R_{i}}(\bar{\theta})+\sum_{k=1}^{M}\nu^{*}_{k}\nabla J_{C_{k}}(\bar{\theta})=0.$
Also, by the complementary slackness of the KKT condition, $\nu^{*}_{k}=0$ for $J_{C_{k}}(\bar{\theta})<d_{k}$. By setting $\bar{\lambda}_{1}=1/(1+\sum_{i=2}^{N}\lambda^{*}_{i})$, $\bar{\lambda}_{i\geq 2}=\lambda^{*}_{i}/(1+\sum_{i=2}^{N}\lambda^{*}_{i})$, and $\bar{\nu}_{k}=\nu^{*}_{k}/(1+\sum_{i=2}^{N}\lambda^{*}_{i})$, the above equation can be rewritten as follows:
$\sum_{i=1}^{N}\bar{\lambda}_{i}\nabla J_{R_{i}}(\bar{\theta})=\sum_{k=1}^{M}\bar{\nu}_{k}\nabla J_{C_{k}}(\bar{\theta}),$ (29)
where $\sum_{i}\bar{\lambda}_{i}=1$, $\bar{\nu}_{k}=0$ for $J_{C_{k}}(\bar{\theta})<d_{k}$. Therefore, the policy $\pi_{\bar{\theta}}$ is constrained-Pareto stationary.
We now investigate the proposition that a policy $\pi_{\bar{\theta}}$ is constrained-Pareto optimal if it is constrained-Pareto stationary, under the assumption that the objectives are concave. Since $\pi_{\bar{\theta}}$ is constrained-Pareto stationary, there exist $\lambda\in\mathbb{R}^{N}_{\geq 0}$ and $\nu\in\mathbb{R}_{\geq 0}^{M}$ satisfying the following:
$\sum_{i=1}^{N}\lambda_{i}\nabla J_{R_{i}}(\bar{\theta})=\sum_{k=1}^{M}\nu_{k}\nabla J_{C_{k}}(\bar{\theta}),$ (30)
where $\sum_{i}\lambda_{i}=1$, $\nu_{k}=0$ for $J_{C_{k}}(\bar{\theta})<d_{k}$. Since $\sum_{i}\lambda_{i}=1$, the value of $\lambda$ in at least one dimension is greater than 00. Without loss of generality, we can refer to the index of such a dimension as $1$. Then, the following is satisfied:
$-\nabla J_{R_{1}}(\bar{\theta})+\sum_{i=2}^{N}\bar{\lambda}_{i}(-\nabla J_{R_{i}}(\bar{\theta}))+\sum_{k=1}^{M}\bar{\nu}_{k}\nabla J_{C_{k}}(\bar{\theta})=0,$ (31)
where $\bar{\lambda}_{i}=\lambda_{i}/\lambda_{1}$ for $k\geq 2$ and $\bar{\nu}_{k}=\nu_{k}/\lambda_{1}$. The above equation can be viewed as the stationarity condition of the KKT conditions for the following optimization problem:
$\mathrm{minimize}_{\theta}-J_{R_{1}}(\theta)\;\textbf{s.t.}\;J_{R_{i}}(\bar{\theta})\leq J_{R_{i}}(\theta)\;\forall i\geq 2,\;J_{C_{k}}(\theta)\leq d_{k}\;\forall k.$ (32)
Since the complementary slackness and the feasibility condition are also satisfied due to (31), all of the KKT conditions are satisfied at $\theta=\bar{\theta}$. Also, the objective and constraints of (32) are convex under the assumptions, so the policy at $\theta=\bar{\theta}$ is the optimal solution of (32). This means that the objectives cannot be improved further within the feasible region, which means the policy is constrained-Pareto optimal. As a result, if a policy is constrained-Pareto stationary, the policy is constrained-Pareto optimal. ∎
Lemma A.1.
If (8) has no solution, the current policy $\pi_{\theta_{\mathrm{old}}}$ is $\epsilon$-CP stationary.
Proof.
Let us assume that the initial policy is feasible, which means $J_{C_{k}}(\theta_{0})\leq d_{k}\;\forall k$. Then, the current policy also satisfies $J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}\;\forall k$ by Theorem 4.2. We can define sets $A:=\{g|\omega_{i}e_{i}\leq g_{i}^{T}g\;\forall i\}$ and $B:=\{g|b_{k}^{T}g+J_{C_{k}}(\theta_{\mathrm{old}})\leq d_{k}-K\epsilon\;\text{for}\;k\in I_{\mathrm{AC}}^{\epsilon}\}$. Equation (8) has no solution means $A\cap B=\emptyset$. Since $A\supset\{g|e_{i}\leq g_{i}^{T}g\;\forall i\}$ and $B\supset\{g|b_{k}^{T}g\leq h_{k}:=\mathrm{min}(d_{k}-K\epsilon-J_{C_{k}},0)\;\text{for}\;k\in I_{\mathrm{AC}}^{\epsilon}\}$,
$\{g|e_{i}\leq g_{i}^{T}g\;\forall i\}\cap\{g|b_{k}^{T}g\leq h_{k}\;\text{for}\;k\in I_{\mathrm{AC}}^{\epsilon}\}=\{g|e_{i}\leq g_{i}^{T}g\;\forall i,\;-h_{k}\leq-b_{k}^{T}g\;\text{for}\;k\in I_{\mathrm{AC}}^{\epsilon}\}=\emptyset.$ (33)
As $e_{i}\geq 0$ and $-h_{k}\geq 0$, the above equation implies that there is no direction to increase $J_{R_{i}}$ and $-J_{C_{k}}$. If we view the CMORL problem as a modified MORL problem, where $-J_{C_{k}}$ is added as objectives, the current policy is Pareto optimal since there is no direction to improve all objectives. According to Theorem 1 in Désidéri (2012), there exist $\lambda\in\mathbb{R}^{N}_{\geq 0}$ and $\nu\in\mathbb{R}^{|I_{\mathrm{AC}}^{\epsilon}|}_{\geq 0}$ such that,
$\sum_{i}\lambda_{i}\nabla J_{R_{i}}(\theta_{\mathrm{old}})-\sum_{k\in I_{\mathrm{AC}}^{\epsilon}}\nu_{k}\nabla J_{C_{k}}(\theta_{\mathrm{old}})=0,$ (34)
where $\sum_{i}\lambda_{i}+\sum_{k}\nu_{k}=1$. According to Lemma A.4 in (Kim et al., 2023), the set $B$ is not empty under the assumptions that the constraints are convex and $\{\theta|J_{C_{k}}(\theta)<d_{k}^{\prime}\;\forall k\}\neq\emptyset$. If $A$ is empty, there exists $\lambda$ such that $\sum_{i}\lambda_{i}\neq 0$. If $A$ is not empty, there exist $\lambda$ and $\nu$ such that $\sum_{i}\lambda_{i}\neq 0$ and $\sum_{k}\nu_{k}\neq 0$ since $A$ and $B$ are not empty. As a result, there exist $\lambda$ such that $\sum_{i}\lambda_{i}$ is positive. By redefining the multipliers as $\lambda^{\prime}=\lambda/\sum_{i}\lambda_{i}$, $\nu^{\prime}_{k\in I_{\mathrm{AC}}^{\epsilon}}=\nu_{k}/\sum_{i}\lambda_{i}$, and $\nu^{\prime}_{k\notin I_{\mathrm{AC}}^{\epsilon}}=0$, the current policy $\pi_{\theta_{\mathrm{old}}}$ satisfies the $\epsilon$-CP stationary condition. ∎
See 4.7
Proof.
The proof of the latter statement can be shown through Lemma A.1, so we only focus on the former statement. Since the objectives for $i\in I_{\omega}$ are monotonically improved according to Theorem 4.2 and are bounded by $R_{\mathrm{max}}/(1-\gamma)$, the objectives will converge to specific values. At these converged values, let us assume that the policy is not $\epsilon$-CP stationary.
Under this assumption, the solution of (8), denoted as $\bar{g}_{\omega}^{\mathrm{ag}}$, must exist according to Theorem 4.2. Furthermore, the gradients of the objectives $g_{i}$ are non-zero for $i\in I_{\omega}$. This is because if one of the gradients of the objectives are zero, the policy parameters satisfy the Pareto stationary condition by setting $\lambda_{i}=1$ and $\lambda_{j\neq i}=\nu_{k}=0$. Since $e_{i}=\epsilon g_{i}^{T}H^{-1}g_{i}\neq 0$ for $i\in I_{\omega}$, $\bar{g}_{\omega}^{\mathrm{ag}}\neq\mathbf{0}$ and $g_{\omega}^{\mathrm{ag}}\neq\mathbf{0}$. Then if we update the policy parameter to $\theta_{\mathrm{new}}=\theta_{\mathrm{old}}+g_{\omega}^{\mathrm{ag}}/2=\beta_{1}\beta_{2}\bar{g}_{\omega}^{\mathrm{ag}}/2$, the following inequality holds:
$J_{R_{i}}(\theta_{\mathrm{new}})-J_{R_{i}}(\theta_{\mathrm{old}})\geq\frac{\beta_{1}\beta_{2}}{2}(e_{i}-\frac{L\beta_{1}\beta_{2}}{4}||\bar{g}_{\omega}^{\mathrm{ag}}||_{2}^{2})>0.$ (35)
This implies that the objective remains non-stationary, which is contrary to the convergence. Due to this contradiction, the policy at the converged values is $\epsilon$-CP stationary. ∎
Appendix B Experimental Details
@@
In this section, we describe the details of each task, including Safety Gymnasium, legged robot locomotion, and Multi-Objective Gymnasium. We then provide a detailed explanation of the performance metrics named hypervolume and sparsity. Finally, we present implementation details, such as network architectures and hyperparameter settings.
B.1 Task Details
B.1.1 Safety Gymnasium
@@
In the Safety Gymnasium (Ji et al., 2023), we utilize the single-agent safe navigation tasks: SafetyPointGoal1-v0 and SafetyCarGoal1-v0, and the multi-agent safe navigation tasks: SafetyPointMultiGoal1-v0 and SafetyCarMultiGoal1-v0. Snapshots of each task are shown in Figure 4.
In the single-agent tasks, the observation space includes velocity, acceleration, and Lidar of the goal and hazards, and the dimensions are 60 and 72 for the point and car robots, respectively. The action space for both robots is two-dimensional. There are eight hazard areas and one goal, and hazard areas are randomly spawned at the beginning of each episode, while goals are randomly placed whenever the agent reaches the current goal. These tasks originally had a single objective: maximizing the number of goals reached. However, in order to have multiple objectives, we have modified them to include a second objective: minimizing energy consumption. Therefore, the reward is two-dimensional, and we use the original definition for the first reward (please refer to Ji et al. (2023)). The second reward is defined as follows:
@@
B.1.2 Legged Robot Locomotion
The legged robot locomotion tasks (Kim et al., 2023) aim to control bipedal or quadrupedal robots so that their velocity matches a randomly sampled command. This command specifies the target linear velocities in the x and y-axis directions
@@ B.1.2 Legged Robot Locomotion
and the target angular velocity in the z-axis direction, denoted by ($v_{x}^{\mathrm{cmd}},v_{y}^{\mathrm{cmd}},\omega_{z}^{\mathrm{cmd}}$). Snapshots of these tasks are shown in Figure 5. The quadrupedal robot has 12 joints and 12 motors, and the bipedal robot has 14 joints and 10 motors. Each robot is operated by a PD controller that follows the target position of the motors, and the target position is given as an action. Hence, the number of motors corresponds to the dimension of action space. In order to provide enough information for stable control, observations include the command, linear and angular velocities of the robot base, and the position and velocity of each joint. As a result, the dimensions of the observation space are 160 for the quadruped and 132 for the biped. Originally, these tasks have a single objective of following a given command and three constraints: 1) maintaining body balance, 2) keeping the height of CoM (center of mass) above a certain level, and 3) adhering to pre-defined foot contact timing. Therefore, we use the original implementation for the observation, action, and cost functions but modify the reward function to have multiple objectives. The modified version has two objectives: 1) reducing the difference between the current velocity and the command and 2) minimizing energy consumption. Then, the reward function is defined as follows:
@@
B.1.3 Multi-Objective Gymnasium
We utilize the Multi-Objective (MO) Gymnasium (Felten et al., 2023) to evaluate the proposed method in MORL tasks which have no constraints. Among several tasks, we use continuous action tasks, and there are four tasks available: mo-hopper-v4, mo-halfcheetah-v4, mo-lunar-lander-continuous-v2, and water-reservoir-v0. mo-mountaincarcontinuous-v0 is also provided as a continuous action task in the MO Gymnasium, but we exclude it since there is an issue in the reward function. Description of the observation space, action space, and the number of objectives for these tasks are summarized in Table 4, and for more details, please refer to Felten et al. (2023).
@@ B.1.3 Multi-Objective Gymnasium
Water-Reservoir $\subseteq\mathbb{R}^{1}$ $\subseteq\mathbb{R}^{1}$ 2 (excess, deficit)
Lunar-Lander $\subseteq\mathbb{R}^{8}$ $\subseteq\mathbb{R}^{2}$ 4 (success, shaping reward, main fuel, side fuel)
B.2 Metric Details
@@
In this section, we explain in detail the performance metrics used in the main text: hypervolume (HV) and normalized sparsity ($\overline{\mathrm{SP}}$). First, we need to prepare an estimated CP front to calculate such metrics. To do this, we pre-sample a fixed number of equally spaced preferences and roll out the trained universal policy for each pre-sampled preference to estimate objective and constraint values. If the constraint values do not exceed the thresholds, we store the objective values, $(J_{R_{1}}(\theta),...,J_{R_{N}}(\theta))$, in a set. Once this process is completed for all preferences, we can construct an estimated CP font by extracting elements from the collected set that are not dominated by any others. Now, we introduce details on each metric.
@@
As mentioned in the main text, sparsity (SP) has a correlation with HV. An increase in HV implies that the elements of the CP front are moving further away from each other, which increases SP. To remove this correlation, we provide a normalized version as defined in (16), and the calculation process is illustrated in Figure 6(c). Given an estimated CP front, $P$, we normalize each estimated objective value so that the minimum value corresponds to zero and the maximum to one. The normalized objectives are denoted as $\overline{J_{R_{i}}}(\theta)$. We then calculate the average of the squares of the distances between elements in this normalized space. This normalized version still produces a large value when a CP front is densely clustered in particular areas while being sparsely distributed overall. Therefore, it maintains the ability to measure sparsity while removing the correlation with HV.
B.3 Implementation Details
B.3.1 MORL to CMORL using the Lagrangian method

### v2

@@ B.1 Proof of Transformation Process
This theorem can be proved by identifying sequences $\nu_{t,i}^{a},\nu_{t,i}^{b},\lambda_{t,k}^{a},\lambda_{t,k}^{b}$ that satisfy all conditions mentioned in Theorem 4.2. By examining (10), $\nu_{t,i}^{a}$, $\nu_{t,i}^{b}$, $\lambda_{t,k}^{a}$, $\lambda_{t,k}^{b}$, and $\alpha_{t}$ can be deduced as follows:
$\nu_{t,i}^{a}=\frac{\nu^{*}_{i}}{\sum_{j=1}^{N}\nu^{*}_{j}},\nu_{t,i}^{b}=0,\lambda_{t,k}^{b}=\frac{\lambda^{*}_{k}}{\sum_{j=1}^{M}\lambda^{*}_{j}},$
$\lambda_{t,k}^{a}=\min(\max(||\bar{g}_{\omega}^{\mathrm{ag}}||,g_{\mathrm{min}}),g_{\mathrm{max}})\cdot\min\left(\frac{\lambda_{k}^{*}}{\epsilon_{t}\sum_{i=1}^{N}\nu_{i}^{*}},\lambda_{\mathrm{max}}\right),$
$\alpha_{t}=\frac{\epsilon_{t}}{\min(\max(||\bar{g}_{\omega}^{\mathrm{ag}}||,g_{\mathrm{min}}),g_{\mathrm{max}})}.$
Substituting the deduced sequences into (10), we can obtain the same formulation as (9). The deduced sequences satisfy that $\sum\nu_{t,i}^{a}=1$, $\sum\lambda_{t,k}^{b}=1$, $\nu_{t,i}^{b}\in\mathbb{R}_{\geq 0}$ is bounded, $0\leq\lambda_{t,k}^{a}\leq g_{\mathrm{max}}\lambda_{\mathrm{max}}$ is also bounded, and $\alpha_{t}\in[\epsilon_{t}/g_{\mathrm{max}},\epsilon_{t}/g_{\mathrm{min}}]$ follows the Robbins-Monro condition. Additionally, since $\lambda_{k}^{*}$ is an optimal point of the dual problem, it satisfies the KKT condition, which results in $\lambda_{k}^{*}=0$ if $J_{C_{k}}(\theta_{t})<d_{k}$. Consequently, $\lambda_{t,k}^{b}(J_{C_{k}}(\theta_{t})-d_{k})\geq 0$. Now, we need to check whether the $\nu_{t,i}^{a}$ converges to a real number. Given that the learning rate follows the Robbins-Monro condition, the policy converges to a specific policy. Since (5) becomes invariant at this point of convergence, the solution of the dual problem is also fixed. Consequently, $\nu_{t,i}^{a}$ converge to a specific value. ∎
Appendix C Experimental Details
@@
In this section, we describe the details of each task, including Safety Gymnasium [18], legged robot locomotion [20], and Multi-Objective Gymnasium [13]. We then provide a detailed explanation of the performance metrics named hypervolume and sparsity. Finally, we present implementation details, such as network architectures and hyperparameter settings.
C.1 Task Details
C.1.1 Safety Gymnasium
@@
In the Safety Gymnasium [18], we utilize the single-agent safe navigation tasks: SafetyPointGoal1-v0 and SafetyCarGoal1-v0, and the multi-agent safe navigation tasks: SafetyPointMultiGoal1-v0 and SafetyCarMultiGoal1-v0. Snapshots of each task are shown in Figure 9.
In the single-agent tasks, the observation space includes velocity, acceleration, and Lidar of the goal and hazards, and the dimensions are 60 and 72 for the point and car robots, respectively. The action space for both robots is two-dimensional. There are eight hazard areas and one goal, and hazard areas are randomly spawned at the beginning of each episode, while goals are randomly placed whenever the agent reaches the current goal. These tasks originally had a single objective: maximizing the number of goals reached. However, in order to have multiple objectives, we have modified them to include a second objective: minimizing energy consumption. Therefore, the reward is two-dimensional, and we use the original definition for the first reward (please refer to Ji et al. [18]). The second reward is defined as follows:
@@
C.1.2 Legged Robot Locomotion
@@
The legged robot locomotion tasks [20] aim to control bipedal or quadrupedal robots so that their velocity matches a randomly sampled command. This command specifies the target linear velocities in the x and y-axis directions and the target angular velocity in the z-axis direction, denoted by ($v_{x}^{\mathrm{cmd}},v_{y}^{\mathrm{cmd}},\omega_{z}^{\mathrm{cmd}}$). Snapshots of these tasks are shown in Figure 10. The quadrupedal robot has 12 joints and 12 motors, and the bipedal robot has 14 joints and 10 motors. Each robot is operated by a PD controller that follows the target position of the motors, and the target position is given as an action. Hence, the number of motors corresponds to the dimension of action space. In order to provide enough information for stable control, observations include the command, linear and angular velocities of the robot base, and the position and velocity of each joint. As a result, the dimensions of the observation space are 160 for the quadruped and 132 for the biped. Originally, these tasks have a single objective of following a given command and three constraints: 1) maintaining body balance, 2) keeping the height of CoM (center of mass) above a certain level, and 3) adhering to pre-defined foot contact timing. Therefore, we use the original implementation for the observation, action, and cost functions but modify the reward function to have multiple objectives. The modified version has two objectives: 1) reducing the difference between the current velocity and the command and 2) minimizing energy consumption. Then, the reward function is defined as follows:
@@
C.1.3 Multi-Objective Gymnasium
@@
Humanoid $\subseteq\mathbb{R}^{376}$ $\subseteq\mathbb{R}^{17}$ 2 (velocity, energy)
Ant $\subseteq\mathbb{R}^{27}$ $\subseteq\mathbb{R}^{8}$ 3 ($x$-velocity, $y$-velocity, energy)
Walker2d $\subseteq\mathbb{R}^{17}$ $\subseteq\mathbb{R}^{6}$ 2 (velocity, energy)
Swimmer $\subseteq\mathbb{R}^{8}$ $\subseteq\mathbb{R}^{2}$ 2 (velocity, energy)
We utilize the Multi-Objective (MO) Gymnasium [13] to evaluate the proposed method in MORL tasks which have no constraints. Among several tasks, we use the MuJoCo tasks with continuous action space, and there are six tasks available: mo-hopper-v4, mo-humanoid-v4, mo-halfcheetah-v4, mo-ant-v4, mo-walker2d-v4, and mo-swimmer-v4. Description of the observation space, action space, and the number of objectives for these tasks are summarized in Table 1, and for more details, please refer to Felten et al. [13].
C.2 Metric Details
@@
In this section, we explain in detail the performance metrics used in the main text: hypervolume (HV) and normalized sparsity ($\overline{\mathrm{SP}}$). First, we need to prepare an estimated CP front to calculate such metrics. To do this, we pre-sample a fixed number of equally spaced preferences and roll out the trained universal policy for each pre-sampled preference to estimate objective and constraint values. If the constraint values do not exceed the thresholds, we store the objective values, $(J_{R_{1}}(\theta),...,J_{R_{N}}(\theta))$, in a set. Once this process is completed for all preferences, we can construct an estimated CP front by extracting elements from the collected set that are not dominated by any others. Now, we introduce details on each metric.
@@
As mentioned in the main text, sparsity (SP) [32] has a correlation with HV is defined as follows:
$\mathrm{SP}(P):=\frac{1}{|P|-1}\sum_{j=1}^{N}\sum_{i=1}^{|P|-1}(\tilde{P}_{j}[i]-\tilde{P}_{j}[i+1])^{2},$ (25)
where $\tilde{P}_{j}:=\mathrm{Sort}(\{p[j]|\forall p\in P\})$. An increase in HV implies that the elements of the CP front are moving further away from each other, which increases SP. To remove this correlation, we provide a normalized version as defined in (14), and the calculation process is illustrated in Figure 11(c). Given an estimated CP front, $P$, we normalize each estimated objective value so that the minimum value corresponds to zero and the maximum to one. The normalized objectives are denoted as $\overline{J_{R_{i}}}(\theta)$. We then calculate the average of the squares of the distances between elements in this normalized space. This normalized version still produces a large value when a CP front is densely clustered in particular areas while being sparsely distributed overall. Therefore, it maintains the ability to measure sparsity while removing the correlation with HV.
C.3 Implementation Details
In this section, several implementation details including hyperparameter settings and network structures are provided. Before that, we report information on the computational resources. In all experiments, we used a PC equipped with an Intel Xeon CPU E5-2680 and an NVIDIA TITAN Xp GPU. The average training time in the single point goal task is presented in Table 2.
Table 2: Training time on the point single goal task.
CAPQL PD-MORL LP3 CoMOGA (ours)
Time 16h 49m 29s 16h 40m 9s 24h 4m 34s 17h 27m 58s
C.3.1 MORL to CMORL using the Lagrangian method